OpenStudy (anonymous):

What is the sum? Please explain answer if possible. 7/5+(-1/6) A. -7/30 B. 6/11 C. 37/30 D. 47/30

OpenStudy (anonymous):

D

OpenStudy (anonymous):

i thought it was b

OpenStudy (anonymous):

1.233333333.. = x 10x = 12.3333333 10x - x = 12.33333 ...... - 1.23333333333..... 9x = 11.1 (all the rest of the threes are subtracted away by a three! x = 11.1/9 x = 111/90 x = 1 21/90 x = 1 7/30

OpenStudy (anonymous):

1 7/30 = 47/30

OpenStudy (anonymous):

hmmm ok if u say so

OpenStudy (anonymous):

look up 7/5 + -1/6, what do you get?

OpenStudy (anonymous):

1.233333333

OpenStudy (anonymous):

then follow my steps and you see that it equals 1 7/30

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

which improper is 37/30 not 47 sorry

OpenStudy (anonymous):

C

OpenStudy (anonymous):

ohhhhh ok i see it

OpenStudy (anonymous):

Your Welcome

OpenStudy (anonymous):

Thanks for trying to help @YK.PAPI

OpenStudy (anonymous):

you welcome